Suppose 2-year Treasury bonds yield 4.9%, while 1-year bonds yield 2.8%. r* is 1.75%, and the maturity risk premium is zero. Use minus sign for any negative expected inflation rate.

  1. Using the expectations theory, what is the yield on a 1-year bond 1 year from now? Calculate the yield using a geometric average. Do not round intermediate calculations. Round your answer to two decimal places.

    b. What is the expected inflation rate in Year 1? Do not round intermediate calculations. Round your answer to two decimal places.

    c. What is the expected inflation rate in Year 2? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Using the expectations theory, what is the yield on a 1-year bond 1 year from now? Calculate the yield using a geometric average.
=1.049^2/1.028-1=7.043%

b. What is the expected inflation rate in Year 1?
=2.8%-1.75%=1.05%

c. What is the expected inflation rate in Year 2?
=(4.9%-1.75%)*2-1.05%=5.250%